OpenStudy (anonymous):

solve the following systems of equations 2x-y-z=-3 2x+2y+3z=2 3x-3y-z=-4

OpenStudy (anonymous):

first set the equations = to 0

OpenStudy (anonymous):

Which equation? I learned like using elimination method or substitution method and im just confused.

OpenStudy (anonymous):

ok which do you want to use?

OpenStudy (anonymous):

we can use the first one @Austin6i6

OpenStudy (owlcoffee):

a system of equation is nothing more complicated than a "n" number of equations with "n" number of variables. And our objetive when solving one, are the variables that solves all the equations at the simultaneusly. So, let's call the equations 1 , 2 and 3: \[1) 2x-y-z=-3\] \[2) 2x+2y+3z=2\] \[3)3x-3y-z=-4\] So, our objective is ti reduce the number of variables and create two other equations with two variables, and make it a 2x2 system of equations. In order to do that, we will take equation 1), multiply it by -1 and add it to the equation 2): \[1)-2x+y+z=3\] \[2)2x+2y+3z=2\] So adding them up, we will create equation 4): \[4) 3y+4z=5\] And we will do it analogically for another pair.But this time, we have to eliminate the same variable we used to create equation 4): Si I will take equation 1) and multiply it by -3 and equation 3) and multiply it by 2: \[1)-6x+3y+3z=18\] \[3)6x-6y-2z=-8\] Now, we will sum them so we can create equation 5): \[5)3y+z=10\] Now that we have made two equations, with only ttwo variables, all you have to do is solve a 2x2 system. Consisting of: \[4) 3y+4z=5\] \[5)3y+z=10\]

OpenStudy (anonymous):

Do you know anything about matrics

OpenStudy (anonymous):

hwo did you get 18? I thought 3 times 3 is 9?

OpenStudy (owlcoffee):

oops, I thought of it as 6, it's 9 yes.

OpenStudy (anonymous):

so the sum woudl acutally be -3y-5z=1?

OpenStudy (anonymous):

brijackson6 do you know anything about matrics?

OpenStudy (anonymous):

no i dont sorry

OpenStudy (anonymous):

ok, because if you did it would make this problem so simple. nvm stick to whats here.